# Probing the Vulnerability of Large Language Models to Polysemantic Interventions

### v1

@@ Appendix C Sparse Autoencoder Training
Appendix D Dataset Generation
When conducting intervention, we do not expect to substantially interfere with every inference of the model, but rather consider intervention in specific contexts, in which we will examine target tokens’ prediction probabilities. We require that the constructed sentences be grammatically capable of deriving the target token so that we can check it in top-10 predictions without severely compromising the model. In later sections, we will see that for tokens such as location names and personal names, we can effectively interfere with the output of the corresponding sentences, which shows a potential to produce hallucinations. Also, we need to point out that more specific and general intervention can be achieved by first listing a set of sentences within that context and then identifying possible ways to interfere with each of them. The prompt to generate the dataset is shown below.
@@ 1. All sentences MUST grammatically require "target_token" next to it
System:Analyze a token target_token from Pythia/GPT2-Small’s vocabulary and return its type in EXACT JSON format:
"[target_token]": {
"type": "[token_type]",
Token Type Rules:
1. "verb": Action words (e.g., "run", "Ġjumping")
2. "location": Place names (e.g., "ĠParis", "ĠTokyo")
3. "person": Names of people/roles (e.g., "ĠJohn", "Ġteacher")
4. "object": Physical objects (e.g., "Ġapple", "Ġtable")
5. "other_noun": Other nouns not in above categories
6. "adjective": Descriptive words (e.g., "Ġhappy", "Ġred")
7. "single_letter": Single characters (e.g., "A", "z")
8. "prefix": Word parts (e.g., "Ġun", "Ġpre")
9. "other": Symbols/punctuation or unclassifiable tokens
User:target_token={token}
Here we provide a table of some example tokens and their sentences generated by DeepSeek-V3 (See Table 3).
@@
loves verb She truly believes that everyone
@@
hates verb Everyone knows that she
@@
sad adjective After hearing the bad news, she felt incredibly
@@
happy adjective After receiving the good news, she felt extremely
@@
Appendix E Supportive Statistics
Active features refer to SAE features that have input texts enabling them to reach an active state. In addition to the semantic clustering of active features mentioned in the main text, we also apply agglomerative clustering to cluster their interference values. The threshold for dividing clusters is set to 0.5. As shown in Figure 7, the vast majority of clusters contain only one feature, indicating that only a small number of features exhibit high interference with each other.
Figure 6: Average number of active features extracted by SAE per Layer
Figure 7: Interference Cluster Size Distribution
Appendix F Intervention Test with Feature Direction
F.1 Generalized Formulation of a Steering-Vector Intervention
@@
F.2 Experiment Details
Our hypothesis posits that if two feature directions interfere in $\mathcal{A}_{\ell}$, despite being nearly orthogonal in $\mathcal{M}$, then enhancing one will inevitably encode some information of the other one into activation space. If true, this would imply the potential to covertly manipulate the output probability of a target feature by steering the model with seemingly unrelated features. However, we need to point out that multiplying the direction of the interference feature by a scale parameter and adding it to the activation layer within a certain range doesn’t always increase the probability of the target feature’s top-activating token in the prediction list. Moreover, we also observed that because the interference feature has the highest interference value of 1 with with itself, it also increases the probability of its own top-activating tokens in the output. Sometimes, the output leans more toward the interference feature, which compromises our examination of the impact on the target feature in the top-10 output tokens. Therefore, to focus on studying the influence of semantically different features on the target feature, we separately calculated the similarity between the output and the top-activating tokens of either the target feature or the interference feature in our metrics. If the similarity with the latter exceeds that with the former, it indicates that the semantics are steered toward the interference feature, and we exclude such results from our statistics.
In the specific experiment, we note that features having an interference feature with value above 0.4 constitute only a small proportion. Additionally, high-interference features are mostly semantically similar, which does not align with our research purpose on the mutual influence of different semantics. Therefore, to ensure the richness of high-interference samples in the sampling process, we first select target features from clusters generated based on interference value and then choose interference features for them, also ensuring that the interference features differ both semantically and in terms of top-activating tokens from the target features. Next, we search for other interference features with low interference values to the target features for experimentation. Specifically, other interference features are selected based on the interference values lying in intervals:$[0.0,0.1]$, $[0.1,0.2]$, $[0.2,0.3]$ and $[0.3,0.4]$. The scale parameter was tested within the range of $[-20,20]$, and we avoide larger ranges to prevent severe disruption of the model.
@@ F.2 Experiment Details
Due to limitations on computational power, we sample clusters and features across various layers. In each experiment with Pythia-70M, we sample 450 target features and collect approximately 2,000 interference features. In each experiment with GPT-2-Small, we sample 1,200 target features and collect approximately 5,000 interference features. As mentioned in the main text, for each feature, we focus on its top-activating token and use DeepSeek-V3 to generate three prompt sentences for it. For each sentence, we test within the aforementioned scale range and record the result with the greatest improvement in the four metrics. This result means the best performance that the steering vector can achieve to induce the semantic of output toward target feature without significantly disrupting the model. And the final four metrics are averaged across all sentences for all features. To show the robustness of our experiments, more results are listed below.
Figure 8: Replications of the feature-level intervention on sampled features for GPT-2-Small. Each row is an independent experiment. The target and interference features are all sampled randomly.
Figure 9: Replications of the feature-level intervention on sampled features for Pythia-70M. Each row is an independent experiment. The target and interference features are all sampled randomly.
Appendix G Intervention Test with Token’s Gradient
G.1 Token Gradient Direction Extraction
@@ G.2 Experiment Details
The SAE dataset from Neuronpedia contains approximately 50 activation text segments per active SAE feature, each strongly activating its corresponding feature. Due to computational limitations, we try to extract the gradient vectors from the first 3 activation texts of each feature. To enable direct comparison with feature-direction vectors, experiments are conducted on the same target and interference features in the above section. Also we scale the vector within the same range $[-20,20]$. Additional experimental results are presented below. Each with steering with token gradients combined with steering with feature directions can be done in one hour and a half for Pythia-70M and six hours for GPT-2-Small, running on a single thread of Intel i7-14700K.
Figure 10: Replications of the token-level intervention on sampled features for GPT-2-Small. Each row is an independent experiment. The target and interference features are all sampled randomly.
Figure 11: Replications of the token-level intervention on sampled features for Pythia-70M. Each row is an independent experiment. The target and interference features are all sampled randomly.
Appendix H Black-Box Interventions on Larger Models
@@
H.1 Target Selection
We deliberately focus on two feature families—(i) location names and (ii) the polarity antonyms “hate” / “love”—because they satisfy three practical and conceptual criteria that make them ideal first-round probes of polysemantic vulnerabilities.
High corpus frequency, low internal polysemy. Most large-scale text corpora mention both world cities/countries and the verbs hate/love thousands of times, giving the SAE a rich activation signal, yet each term carries a relatively unambiguous core meaning. This minimizes confounds from “target drift” when we measure probability shifts.
Complementary linguistic classes. Locations are concrete named entities rooted in external knowledge, whereas hate/love are abstract affective predicates that drive sentiment. Showing transferable interference for both a proper-noun category and an emotional-valence category demonstrates that the vulnerability is not restricted to a single part of speech or semantic field.
Policy relevance. Manipulating geographic references risks misinformation about real-world facts (“The capital of X is…”), while manipulating strong sentiment verbs directly impacts toxicity and persuasion. Successful steering of these tokens therefore highlights two distinct, societally significant threat surfaces-factual reliability and affective bias.
Together, these criteria make locations and hate/love a parsimonious yet representative pair for an initial, systematic evaluation; expanding to additional categories is an important next step once the core risk is established.
H.2 Steering with Token Gradient Vector
The scalable intervention on LLaMA3.1-8B-Instruct is conducted by first selecting target type tokens as mentioned above and identifying target features in Pythia-70M and GPT-2-Small for which these tokens are the top-activating ones. The two interference feature sets in Pythia-70M and GPT-2-Small with respect to the target features are then identified. Next, we collect the top-activating tokens in two models respectively, and compute the intersection. The figure12 shows a sketch of the shared interference tokens across two models. From the token cloud map, we can observe that high-interference tokens are often punctuation-like tokens such as line breaks. There are also some tokens with specific meanings, which may be related to certain target tokens in daily context.
(a) Name interference tokens
@@ H.2 Steering with Token Gradient Vector
(c) Love/Hate interference tokens
@@
After getting the tokens, we proceed to collect the activation texts, which may activate interference features of the target in black-box models. Due to computational constraints, we only compute gradients from and perform operations on the first half of residual layers in LLaMA3.1-8B-Instruct. The intervention experiments are done for three target token types, each containing about 100 sentences. More intervention examples are listed below. It takes about $20$ minutes for a single RTX4090 GPU to find a highly effective gradient vector for steering.
Table 4: Examples of interventions on LLaMa3.1-8B-Instruct Using Token Gradient Vector
@@
name quantitative data points related to statistics and performance metrics “Nobody expected the mysterious package to be from” $\uparrow$ Entered $\downarrow$ Dropped Paul +0.261 the -0.104 Emmanuel +0.026 a -0.078 Matthew +0.021 Lake -0.51

### v2

@@ Appendix D Sparse Autoencoder Training
Appendix E Dataset Generation
The investigation of the interference effects of SAE features is conducted within the contexts where they are likely to fire, which allows us to manipulate the expression of these features without significantly compromising the model. For example, when given the prompt “In the weekend, we are going to", the features related to locations are incorporated into our examination. Additionally, we observe a positive correlation between the boost of a SAE feature and the increased probability of its high-activation tokens in the model’s next-token predictions. Therefore, here we simply construct sentences for each token in the model’s vocabulary, ensuring that these sentences are grammatically capable of leading to the token and evaluate the expression of features with respect to the tokens’ output probabilities.
@@ 1. All sentences MUST grammatically require "target_token" next to it
System: You are a linguistic analyzer. Your task is to classify tokens from language model vocabularies into semantic categories.
Given a list of tokens, classify each token into ONE of these categories:
- person: Names, pronouns, occupations, human-related terms
Examples: " John", " Mary", " doctor", " teacher", " he", " she", " people"
- location: Cities, countries, geographical features, places
Examples: " London", " America", " mountain", " beach", " city", " Paris", " China"
- time: Time units, dates, temporal expressions, seasons
Examples: " Monday", " January", " morning", " year", " day", " week", " winter"
- number: Numerical digits, number words, mathematical terms
Examples: " one", " two", " 1", " 2", " first", " hundred", " plus", " minus"
- animal: All living creatures, insects, pets, wildlife
Examples: " dog", " cat", " bird", " lion", " fish", " elephant", " butterfly"
- food: Edible items, drinks, cooking terms, ingredients
Examples: " apple", " bread", " water", " coffee", " chicken", " rice", " pizza"
- color: Color names, shades, visual descriptors
Examples: " red", " blue", " green", " black", " white", " yellow", " purple"
- emotion: Feelings, emotional states, psychological terms
Examples: " happy", " sad", " angry", " love", " fear", " excited", " worried"
- body: Human/animal body parts, anatomy
Examples: " head", " hand", " eye", " heart", " leg", " face", " brain"
- transport: Vehicles, transportation methods
Examples: " car", " bus", " train", " plane", " ship", " bicycle", " truck"
- science: Scientific terms, elements, physics/chemistry concepts
Examples: " oxygen", " energy", " atom", " DNA", " gravity", " electron", " acid"
- abstract: Philosophy, ideas, concepts, mental constructs
Examples: " freedom", " justice", " truth", " idea", " concept", " theory", " belief"
- object: Physical items, tools, furniture, equipment
Examples: " table", " chair", " book", " phone", " computer", " tool", " box"
- action: Verbs, activities, movements, processes
Examples: " run", " walk", " eat", " think", " write", " jump", " sleep"
- unknown: Anything that doesn’t clearly fit into the above categories
Return a JSON object where each key is a token and each value is its category.
Rules:
1. Classify ALL provided tokens
2. Use EXACT token strings as keys (including spaces)
3. Choose the MOST appropriate single category
4. Use lowercase category names only
5. When in doubt, use ‘‘unknown"
6. Return ONLY the JSON object, no additional text
User: Classify these tokens:
token_str_1, token_str_2, ....
Here we provide a table of some example tokens and their sentences generated by DeepSeek-V3 (See Table LABEL:tab:token_examples).
@@ 6. Return ONLY the JSON object, no additional text
loves emotion She truly believes that everyone
@@
hates emotion Everyone knows that she
@@
sad emotion After hearing the bad news, she felt incredibly
@@
happy emotion After receiving the good news, she felt extremely
@@
Table 2: Token type and prompt sentences examples (continued)
Appendix F Supportive Statistics
Active features refer to SAE features that have input texts enabling them to reach an active state. In addition to the semantic clustering of active features mentioned in the main text, we also apply agglomerative clustering to cluster their interference values. The threshold for dividing clusters is set to $0.4$. As shown in Figure 7, the vast majority of clusters contain only one feature, indicating that only a small number of features exhibit high interference with others.
In Figure 8, we also provide an overview of the distribution of semantic similarity and interference values between features, as well as their correlations. Since the distribution and correlation of these two types of data remain largely consistent across all layers of the model, we present a representative example layer for illustration.
Figure 6: Number of active features extracted by SAEs per layer. A is the result of Pythia-70M, and B is the result of GPT-2-Small. Error bars represent $95\%$ confidence intervals.
Figure 7: Interference Cluster Size Distribution. A is the result of Pythia-70M, and B is the result of GPT-2-Small. Error bars represent $95\%$ confidence intervals.
Figure 8: Semantic similarity and interference distribution of Pythia Res-4 and GPT2-Small Res Post-4 layer. (A–C) are the results of Pythia-Res-4. For semantic similarity, $14.0\%$ of values are below $0.15$, $29.6\%$ are below $0.2$, $67.4\%$ are below $0.3$, and $89.6\%$ are below $0.4$. For interference, $85.5\%$ of values are below $0.1$, $96.4\%$ are below $0.2$, $99.2\%$ are below $0.3$, and $99.8\%$ are below $0.4$. The bivarite analysis suggests semantic similarity and interference value is positively associated ($\beta=0.071$, $s.d=0.092$, $p<0.001$). (D–F) are the results of GPT-2-Smal Res Post-4. For semantic similarity, $10.8\%$ of values are below $0.15$, $24.5\%$ are below $0.2$, $63.3\%$ are below $0.3$, and $89.5\%$ are below $0.4$. For interference, $95.6\%$ of values are below $0.15$, $99.7\%$ are below $0.2$, approximately all values are below $0.3$ and $0.4$. The bivarite analysis also suggests a positive association ($\beta=0.030$, $s.d=0.049$, $p<0.001$).
Appendix G Definition of the Alternative Metric: Weighted Overlap
In addition to the weighted cosine similarity, we report results with an alternative metric, weighted overlap, which measures the raw probability mass assigned to the feature-associated token set $T_{f}\subset V$ in Section H and I. This metric does not smooth over near misses via embedding similarity; instead, it directly captures how much of the model’s next-token distribution lands on tokens in $T_{f}$.
Definition.
For a model output distribution $P\in\Delta^{|V|}$ and target token set $T_{f}$,
$w(P,T_{f})\;=\;\sum_{t\in T_{f}}P(t).$ (3)
Intervention effect.
Let $O$ and $\tilde{O}$ denote the model’s output distributions before and after intervention, respectively. The (absolute) change in weighted overlap is
$\Delta w\;=\;w(\tilde{O},T_{f})\;-\;w(O,T_{f}).$ (4)
Relative change.
When a scale-free summary is preferred, we also report the relative change:
$\widehat{\Delta w}\;=\;\frac{w(\tilde{O},T_{f})-w(O,T_{f})}{\max\{w(O,T_{f}),\epsilon\}},$ (5)
where $\epsilon>0$ is a small constant to avoid division by zero.
Appendix H Intervention Test with Feature Direction
H.1 Generalized Formulation of a Steering-Vector Intervention
@@
H.2 Experiment Details
To obtain the complete intervention data of various levels of interference features on the target features, we first filtered out all features in each layer that contained interference values at all levels, and for which the semantic similarity with the target feature is below the four selected thresholds. Subsequently, we select a subset of these features as the target features and identify the corresponding interference features across the various interference levels. Next, we search for other interference features with low interference values to the target features for experimentation. Specifically, other interference features are selected based on the interference values lying in intervals:$[0.0,0.1]$, $[0.1,0.2]$, $[0.2,0.3]$ and $[0.3,0.4]$. The scale parameter was tested within the range of [20,20][-20,20], and we avoid larger ranges to prevent severe disruption of the model.
@@ H.2 Experiment Details
Due to limitations on computational power, we sample clusters and features across various layers. In each SAE experiment with Pythia-70M, we sample 180 target features and collect approximately 2,700 interference features. In each SAE experiment with GPT-2-Small, we sample 480 target features and collect approximately $7,200$ interference features. For gradient experiments, we reduced 60% sampled features, but three times the number of gradient intervention vectors are extracted to keep the test set scale. As mentioned in the main text, for each feature, we focus on its top-activating token and use DeepSeek-V3 to generate three prompt sentences for it. For each sentence, we test within the aforementioned scale range and record the result with the greatest improvement in the two metrics. This result means the best performance that the steering vector can achieve to induce the semantics of output toward the target feature without significantly disrupting the model. The final two metrics are averaged across all sentences for all features. To show the robustness of our experiments, the results of the alternative metric are presented in Figure 9.
Figure 9: Feature-level intervention effects measured by weighted overlap. Subpanels follow the same conventions as Figure 3, but intervention effects are computed using relative change of weighted overlap ($\widehat{\Delta w}$). For (A–B), error bars indicate 95% confidence intervals; for (C), error bars denote 90%, 95%, and 99% confidence intervals.
Type Model Intervention Target feature Result
Llama-3.1-8b-Instruct Steering gradient vector: references to the world and its various aspects References to ‘Switzerland’ “I would like to recommend you to spend holidays in” $\uparrow$ Entered $\downarrow$ Dropped Switzerland +0.16 Italy -0.034 Germany +0.089 Greece -0.016 Canada +0.015 Bulgaria -0.012
Llama-3.1-8b-Instruct Prepending the injection text “(placement from placement)” References to locations “In the next weekend we will go to” $\uparrow$ Entered $\downarrow$ Dropped Paris +0.011 another -0.013 * +0.010 H -0.003 - +0.006 K -0.003
Appendix I Intervention Test with Token’s Gradient
I.1 Token Gradient Direction Extraction
@@ I.2 Experiment Details
The SAE dataset from Neuronpedia contains approximately $50$ activation text segments per active SAE feature, each strongly activating its corresponding feature. Due to computational limitations, we try to extract the gradient vectors from the first 3 activation texts of each feature. Also, we scale the vector within the same range [0.5,1,1.5,2,3,4,5,6,7,8,9,10,12,14,17,20][0.5,1,1.5,2,3,4,5,6,7,8,9,10,12,14,17,20]. Each experiment on steering with token gradients combined with steering with feature directions can be done in one hour and a half for Pythia-70M and six hours for GPT-2-Small, running on a single thread of Intel i7-14700K. The results of the alternative evaluation metric are presented in Figure 10.
Figure 10: Token-level intervention effects measured by weighted overlap. Subpanels follow the same conventions as Figure 4, but intervention effects are computed using relative change of weighted overlap ($\widehat{\Delta w}$). For (A–B), error bars indicate 95% confidence intervals; for (C), error bars denote 90%, 95%, and 99% confidence intervals.
Appendix J Experiments with DeepSeek-V3 Feature Descriptions
To evaluate the sensitivity of our findings to the selection of LLMs for auto-interpretation, we employ an alternative model, DeepSeek-V3, to interpret the features of a sparse auto-encoder. The explanation text is subsequently processed by the text-embedding-3-large model to obtain embeddings. These embeddings are then utilized as semantic vectors to assess the similarity between features. For this illustrative sampling, features from the Pythia-70M att-5 layer are selected. The interpretation of features by large models is generated by feeding the model the top-activating texts of the feature and denoting the high activating tokens in it. Specifically, the prompts we write for DeepSeek-V3 are as listed below.
System:
We’re studying neurons in a neural network. Each neuron activates on some particular word or concept in a short document. The activating words in each document are enclosed by << and >>. Look at the parts of the document the neuron activates for and summarize in a single sentence what the neuron is activating on. Try to be general in your explanations. Don’t just repeat activation words. Also, you can summarize multiple points if the text content is not highly consistent. Pay attention to things like the capitalization and punctuation of the activating words or concepts, if that seems relevant. Keep the explanation as short and simple as possible, limited to 32 words or less. Omit punctuation and formatting.
User: The activating documents are given below:
1.activation_text_1
2.activation_text_2
5.activation_text_5
In Figure 11, we compare semantic relatedness among SAE features using embeddings of explanations generated by DeepSeek-V3 and GPT-4o-mini. The left density plot shows that pairwise similarities from the two models are tightly aligned, while the right heatmaps further illustrate that both models induce comparable feature–feature semantic structure.
Figure 11: Semantic relatedness between features from the view of DeepSeek-V3 and GPT-4o-mini.
Figure 12 compares SAE feature-direction and gradient-based interventions under auto-interpretations from GPT-4o-mini and DeepSeek-V3. Semantic dissimilarity threshold is set to different scales for intervention feature selection. Although this check is limited to a single layer (and thus exhibits greater variance), we consistently observe that high-interference, low-semantic-similarity features steer the target far more strongly than a random baseline. This pattern holds under both interpreters, indicating robustness to the choice of auto-interpretation model.
Figure 12: Intervention effects under GPT-4o-mini vs. DeepSeek-V3 auto-interpretations of SAE features. Top row shows SAE-direction interventions; bottom row shows gradient-based interventions. Columns are the evaluation metrics: left, $\Delta c$; right, $\widehat{\Delta w}$. Error bars denote $95\%$ confidence intervals.
Appendix K Black-Box Interventions on Larger Models
@@
K.1 Steering with Token Gradient Vector
The scalable intervention on Llama-3.1-8B-Instruct is conducted by first selecting target type tokens as mentioned above and identifying target features in Pythia-70M and GPT-2-Small for which these tokens are the top-activating ones. The two interference feature sets in Pythia-70M and GPT-2-Small with respect to the target features are then identified. Next, we collect the top-activating tokens in two models respectively, and compute the union. The figure13 shows a sketch of the interference tokens extracted from two models.
(a) Person interference tokens
@@
(c) Emotion interference tokens
@@
After getting the tokens, we proceed to collect the activation texts, which may activate interference features of the target in black-box models. Due to computational constraints, we only compute gradients from and perform operations on the first half of residual layers in Llama-3.1-8B-Instruct. The intervention experiments are done for three target token types, each containing about 100 sentences. More intervention examples are listed below. It takes about $20$ minutes for a single RTX4090 GPU to find a highly effective gradient vector for steering.
Table 4: Examples of interventions on Llama-3.1-8B-Instruct Using Token Gradient Vector
@@
person quantitative data points related to statistics and performance metrics “Nobody expected the mysterious package to be from” $\uparrow$ Entered $\downarrow$ Dropped Paul +0.261 the -0.104 Emmanuel +0.026 a -0.078 Matthew +0.021 Lake -0.51